Grade 1210 grade science

The radius of curvature of a spherical mirror is 20 cm. What is its fucul length?

Pawan Prajapati
5 Years agoGrade 12

2 Answers

Harshit Singh
5 Years ago
Dear Student

R = 20cm
R = 2f
f = R/2
f = 20/2
f = 10cm

Thanks
KAPADIA DEV
5 Years ago

Radius of curvature (R) = 20 cm

Radius of curvature of a spherical mirror = 2 — Focal length (f)

R= 2f

f= R/2

f = 20 / 2

f = 10

Hence, the focal length of the given spherical mirror is 10 cm